HUBBLE SPACE TELESCOPE DAILY REPORT #2187 PERIOD COVERED: 0000Z (UTC) 08/13/98 - 0000Z (UTC) 08/114/98 Daily Status Report as of 226/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 7622 (WF/PC-2 Cycle 7 Internal Monitor) The WF/PC-2 was used to make a set of Cycle 7 internal monitoring calibration measurements. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. A bias calibration and an internal flat calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed NIC/1/2/3 7831 (NICMOS Study of Nova Shells) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the nova shock shell around nova QV-Vul. The angular expansion rates and spatial distribution of the ejecta in both the continuum and emission lines of ionized gas will be determined. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed FGS 7681 (FGS #3 Red Transfer Function Reference Calibration) The Fine Guidance Sensors (FGS #3, the astrometer) were used to make observations of SAO185689 in order to perform a red transfer function reference calibration through the PUPIL, CLEAR and NEUTRAL DENSITY filters at the center of the FGS #3 field of view (FOV). The observations were completed as planned, and no anomalies were reported. 1.6 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.7 Completed Four Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations take in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.8 Completed NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.9 Completed NIC/2/3 7453 (The SBF Hubble Diagram) The Near Infrared Camera (Camera Two and Three) was used to make deep NICMOS observations of galactic cluster Abell-3656 to determine its distance using the surface brightness fluctuation {SBF} method. HST is essential for its high spatial resolution which makes it possible to detect and remove globular clusters and background galaxies that contaminate the SBF measurement. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed NIC/1/2/3 7826 (Near-IR Photometry of M-31's Metal-Rich Globular Clusters) The Near Infrared Camera (Cameras One, Two and Three) were used to make observations of the metal-rich globular cluster G174 in galaxy M-31-G174. The high metallicities of the program clusters will make these observations of particular interest for furthering our understanding of the evolution of metal rich stars, and for the population synthesis models meant to study the stellar content of distant, unresolvable bulges and elliptical galaxies. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed WF/PC-2 8059 (POMS Test Proposal: WF-2 Targeted Parallel Archive Proposal) The WF/PC-2 was used to make observations in parallel with the prime science observations. The parallel opportunities available with WF/PC-2 in the neighborhood of bright galaxies are treated in a slightly different way from the normal pure parallels. Local group galaxies offer the opportunity for a closer look at young stellar populations. The observations were completed as planned, and no anomalies were reported. 1.12 Completed STIS/CCD 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a non-scripted parallel POMS test. The observations were completed as scheduled, and no problems were reported. 1.13 Completed WF/PC-2/STIS/CCD/NIC/1/2/3 7307 (The Formation and Evolution of Rich Star Clusters in the LMC) The WF/PC-2, the Space Telescope Imaging Spectrograph (CCD), and the Near Infrared Camera (Cameras One, Two and Three) were used to make observations of the rich star clusters which are located within the Large Magellanic Cloud in order to study their formation and evolution. These observations will help us understand the origin and evolution of rich star clusters in our own galaxy. The observations were completed as scheduled, and no anomalies were reported. 1.14 Completed STIS/MA1 7298 (A Snapshot Survey for UV bright Quasars at redshift below 3) The Space Telescope Imaging Spectrograph (MA1) was used to make ultraviolet snapshot observations of the bright quasar CS0609 which has a redshift (Z) below 3.0. The goal is to find quasars bright enough in the ultraviolet for spectroscopic follow-up in the range 900 -- 1200 Angstroms. A wave length calibration image was also taken. The observation were completed as scheduled, and no anomalies were reported. 1.15 Completed FOC/96 7923 (FOC F/96 Turnon) The Faint Object Camera (f/96) was used to take an internal flat and a dark to ensure that the 30-day limit for the FOC is not exceeded. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed NIC/1/2/3 7875 (Searching for Old Stellar Populations in Candidate Proto-Spheroidals) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations searching for old stellar populations in candidate proto-spheroid SA57-3605. We have compelling evidence from WF/PC-2 images that these galaxies have optical colors gradients. The near-infrared is necessary, however, to reveal color gradients large enough to separate young and old stellar populations. Only NICMOS imaging provides enough spatial resolution and depth to resolve these gradients. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 9 Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 20 Successful : 20 2.3 Operations Notes: No HSTARs were reported. The IP Testing for TDRSS-West was successfully completed at 225/1401Z through 1452Z (OR 13952). During the event, the STOCC received 4 kbps AN format telemetry and transmitted 1 NOOP and 11 FIVEBLCK procedures in COUNTCHECK Mode and 1 NOOP and 15 procedures in the NOVERIFY Mode for a total of 3436 commands with no problems. The PRS Release 11.3 Burn-in was started, and will continue though the week-end. The NICMOS Suspend RTCS (Multiple TPG reset commanding) patch was uplinked at 226/1145Z. This included single word patch NICMOS Suspend sequence from Sequence 8 through Sequence 25. NASCOM VDS Y2K Compliance was started and continues through the weekend. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH